A steam power plant is intended to be used for heating as well as power generation. The net power required from this installation is 5 MW and the heating power is 12 MW. While the steam produced at 4 MPa, 440 degrees Celsius in the boiler expands in the isentropic turbine, the steam to be used in heating is drawn from the turbine at 140 kPa and it is condensed in the heating installation and mixed with water from the condenser with a pressure reducer after cooling down to 40 degrees Celsius. In addition, intermediate steam is taken at 0.4 MPa to be sent to ten mixed mixtures. The pumps are isentropic and the condenser temperature is 40 degrees Celsius.

a) Calculate the flow rate of the steam used for heating.

b) Find the flow rate of the steam and intermediate steam produced in the boiler.

C) Calculate the heat given in the boiler and the heat taken from the condenser at once.
Answer #1

Ampa, 440°C boiler Turbine z? NE 2 3 о•4мра, Pa 14okran I js =m-16,76) & wytā ondenses X50 Trap 40°C 1 A S T-S diagram.heated T,-ho°, P, - км Р Supe vapor. А, - 331obf19 : Si= 6.910 kJ l kyk P2=0rumpa, S = 6.410 kt (kgr Superheated vapor. Т.: +Ps= P3=1 houpa, *5=0 ts=109°C hs=458 kilky haya hotu6(P7-86) = 168 + 1.0078 X163( two-T38 ) hy 168.39568 kJ / kg P8=0.4mpa, X4 a) j chy-hs) = 12x103 %₂/2560-458) = 12x103 jea - sito88 kg/see b) Y + G + z = in Nef power, wz-lipz-wpg=5x10 [is, (home)+s• 39sЬ8 во 19 S7o5, +16ой, - O» 2956° 0,-де. 3. 888 4. – 22.19ь- 3.886 ч, 211** s46-12-у, - о•39 sex s «тоғf +155 тый, chi8-9• - 2, 3, ha. - 1.71 +5, 7o88 + 0.047 in = 6.9328 kg/sen c) aina nchinha) -6.932/( 3310 ~ 608.888) = 18 726.269 kW Qin = 18,7